# EAMA : Entity-Aware Multimodal Alignment Based Approach for News Image Captioning

### v4

@@ 6. Conclusion
Figure 4. An example of our prompts for Entity-Aware Sentence Selection, Entity Selection and News Image Captioning tasks. We label entities in the caption in green. We label entities in the input context but not in the caption in blue.
@@ Appendix B Implementation details
We restrict the max length of multimodal input tokens to 1024 and the max length of output new tokens to 50 on most experiments. We set the max length of multimodal input tokens to 2048 while taking the full news article as textual input for the specific experiment in Section 5.4. Following previous researches, the traditional setting to obtain textual input on the GoodNews dataset involves extracting the first 500 words of the entire news article. Conversely, on the NYTimes800k dataset, the traditional textual input(as loc_context) entails obtaining words around the image, as the NYTimes800k dataset provides information about the image position within the entire article. Our method extracts sentences which contain entities relevant to the given news image from the entire news article first, then applies these sentences to supplement the textual input context. We set a maximum limitation of 600 words for our supplemented textual input context. In our Balance Training Strategy setting, we organize the data into mini-groups, each containing 2 instances of the News Image Captioning task, 1 set of instances of the Entity-Aware Sentence Selection task, and 1 instance of the Entity Selection task. Following previous researches, we employ SpaCy to recognize entities during our research.
In the OSFT setting, we utilize the official scripts for InstructBLIP, MiniGPT-v2, and LLaVA-v1.5. Since there are no official scripts available for fully finetuning InstructBLIP, we re-implement InstructBLIP to enable full finetuning with deepspeed. All training hyper-parameters follow their default configurations: the initial learning rates for InstructBLIP and LLaVA-v1.5 are set to $52e-$ 5, while the initial learning rate for MiniGPT-v2 is set to $51e-$ 5. During inference, we use a beam size of 5 in MLLMs. Mixed precision training with bp16 is employed to accelerate the training process. Our model takes approximately 2-3 days to train on four Nvidia A40 GPUs.
Appendix C Analysis of Entity Generation
@@
        EAMA         50.06         10.25
Table 5. The Recall of Entities in and out of textual input context on NYTimes800k dataset. All methods except EAMA take loc_context as textual input, EAMA takes supplemented textual input. We cite the results of Tell and JoGANIC from previous researches.
In-Out Textual Input Entity Generation According to prior researches, evaluating the generation of entities that appear in both the textual input context and captions can provide better insight into the models’ ability on handling entity information. Therefore, we further analyze the Recall of entities in and outside of the textual input context.
@@
All results are listed in Table 5, InstructBLIP obtain higher Recall especially for entities of caption occurring in the textual input context. Our method obtains the highest Recall of entities appearing in textual input context, which is consistency with the our previous results.
@@
         EAMA          29.33          27.54
Table 6. The Precision and Recall of Entities which are out of train dataset in our approach: EAMA, on NYTimes800k dataset.
Out-of-Train Entity Generation To assess our method’s capability to predict entities not present in the training data, we examine captions in the NYTimes800k test dataset, focusing on entities that do not appear in the training data. The entire test set comprises 70,002 entities in the reference captions, with 13,481 of them not appearing in the textual input contexts or captions of the training data. Our approach, EAMA, successfully predicts 3,713 entities in the generated captions, resulting in a Recall of 27.54. The Precision of the out-of-train entities is 29.33, as shown in Table 6. Both the Recall and Precision metrics are comparable to the results for all entities presented in Section 5.2, demonstrating the generalization ability of EAMA for generating out-of-train entities.
Figure 5. An example of news image caption generation. We present the news image, news article and the news image caption of one sample from NYTimes800k dataset. The news image captions are respectively, generated in: (1) InstructBLIP in zero-shot setting. (2) LLaVA-v1.5 in zero-shot setting. (3) MiniGPT-v2 in zero-shot setting. (4) GPT-4V in zero-shot setting. (5) InstructBLIP (OSFT) on NYTimes800k. (6) LLaVA-v1.5 (OSFT) on NYTimes800k. (7) MiniGPT-v2 (OSFT) on NYTimes800k. (8) Our method, EAMA, on NYTimes800k. (9) is the ground-truth caption. We highlight Entities occurred in reference in red.
Figure 6. An example of news image caption generation. We present the news image, news article and the news image caption of one sample from NYTimes800k dataset. The news image captions are respectively, generated in: (1) InstructBLIP in zero-shot setting. (2) LLaVA-v1.5 in zero-shot setting. (3) MiniGPT-v2 in zero-shot setting. (4) GPT-4V in zero-shot setting. (5) InstructBLIP (OSFT) on NYTimes800k. (6) LLaVA-v1.5 (OSFT) on NYTimes800k. (7) MiniGPT-v2 (OSFT) on NYTimes800k. (8) Our method, EAMA, on NYTimes800k. (9) is the ground-truth caption. We highlight Entities occurred in reference in red.
Appendix D Examples
We show two examples of case study in figure 5 and 6 of captions generated by different MLLMs on our research. Our method achieves the best generation on entities and captions on both examples.

### v5

@@ Appendix A Training and Inference Prompts
Dataset MLLM Setting BLEU-4 METEOR ROUGE CIDEr Named entities
P R
GoodNews LLaVA-v1.5 Zero_Shot 2.63 8.72 14.03 15.85 17.19 10.96
OSFT 8.48 12.37 25.22 78.03 25.69 24.41
MiniGPT-v2 Zero_Shot 1.81 7.62 11.81 9.77 15.19 10.17
OSFT 8.05 12.22 22.78 62.42 21.32 23.62
InstructBLIP Zero_Shot 2.52 6.68 10.78 14.82 13.34 8.78
OSFT 9.53 13.54 25.61 78.03 25.89 27.33
NYTimes800k LLaVA-v1.5 Zero_Shot 2.65 8.75 13.53 14.66 17.24 11.82
OSFT 7.94 11.67 23.48 65.28 25.92 25.01
MiniGPT-v2 Zero_Shot 1.85 7.69 11.56 10.21 14.15 11.41
OSFT 7.38 11.32 21.27 53.30 21.25 23.89
InstructBLIP Zero_Shot 2.52 7.41 11.01 12.74 12.79 11.21
OSFT 10.05 13.63 25.45 75.95 27.39 30.37
Table 3: Performance of MLLMs in Zero_Shot and Official supported Supervised Finetune (OSFT) setting. Following the OSFT settings, we train the V-L Connector in MiniGPT-v2 and InstructBLIP, and fully train the V-L Connector together with LLM in LLaVA-v1.5. LLaVA-v1.5, MiniGPT-v2 and InstructBLIP are implemented with 7B checkpoints.
@@ Appendix B Implementation details
We restrict the max length of multimodal input tokens to 1024 and the max length of output new tokens to 50 on most experiments. We set the max length of multimodal input tokens to 2048 while taking the full news article as textual input for the specific experiment in Section 5.4. Following previous research(Tran et al., 2020), the traditional setting to obtain textual input on the GoodNews dataset involves extracting the first 500 words of the entire news article. Conversely, on the NYTimes800k dataset, the traditional textual input(as loc_context) entails obtaining words around the image, as the NYTimes800k dataset provides information about the image position within the entire article. Our method extracts sentences which contain entities relevant to the given news image from the entire news article first, then applies these sentences to supplement the textual input context. We set a maximum limitation of 600 words during this stage. Our method will then extract the related entities from the news article and combine them with prompt such as "The possible related entities are:" into textual input context. In Entity-Aware Multimodal Alignment stage, we organize the data into mini-groups, each containing 2 instances of the News Image Captioning task, 1 set of instances of the Entity-Aware Sentence Selection task, and 1 instance of the Entity Selection task. Following previous researches, we employ SpaCy to recognize entities during our research.
In the OSFT setting, we utilize the official scripts for InstructBLIP, MiniGPT-v2, and LLaVA-v1.5. Since there are no official scripts available for fully finetuning InstructBLIP, we re-implement InstructBLIP to enable fully finetuning with deepspeed as Base in our experiments.
When aligning our model with the two Entity-Aware alignment tasks, the final loss is the weighted sum of loss in there tasks. We set the weight of $\mathcal{L}^{\mathrm{SENT}}_{\mathrm{LM}}$, $\mathcal{L}^{\mathrm{ENT}}_{\mathrm{LM}}$, and $\mathcal{L}_{\mathrm{LM}}^{\mathrm{CAP}}$ as follows: 0.5, 0.25, and 1 respectively for the GoodNews dataset, and 0.25, 0.75, and 1 respectively for the NYTimes800k dataset. These values are determined based on their performance on the validation dataset.
Appendix C Performance of Different MLLMs
We present the whole evaluation results of different MLLMs in Table 3. The results correspond with our observation that MLLMs perform badly under zero_shot settings and simple finetuning yields promising results, revealing the potential of MLLMs on News Image Captioning task.
Method Textual Input Textual Input BLEU-4 METEOR ROUGE CIDEr Named entities
NIC Training NIC Inference P R
Base Origin Origin 10.88 14.09 26.60 82.70 29.22 31.32
Base Full Full 10.76 13.84 26.18 83.17 27.71 30.03
Base Origin Origin(Longer) 10.60 13.80 26.28 81.75 28.95 30.57
Base Origin Oracle(SENT) 11.24 14.50 27.70 87.56 33.42 34.80
Base Origin Oracle(ENT) 11.63 14.76 27.90 90.44 32.04 34.24
Base Origin Oracle(SENT+ENT) 11.89 15.07 28.79 94.27 35.71 37.10
Align Origin Origin 10.80 14.08 26.80 84.45 29.15 31.45
Align Origin Origin(Longer) 10.75 14.00 26.70 83.50 29.25 31.16
EAMA Origin Supplemented 11.03 14.22 27.15 87.00 29.79 32.24
Table 4: Analysis of Textual Input in News Image Captioning (NIC) on NYTimes800k dataset. Base indicates fully finetuning InstructBLIP on the NIC task. Align means align InstructBLIP with all three alignment tasks. Origin denotes the traditional textual input of NIC. Full denotes the entire news article. Supplemented denotes fusing the extracted sentences and entities with Origin under a limited length. Origin(Longer) denotes obtaining textual input of the same maximum length as Supplemented but using traditional setting. Oracle(SENT) denotes using the sentences containing the same entities as captions and Oracle(ENT) denotes using entities exactly both appearing in the new article and caption. Oracle(SENT+ENT) denotes using the oracle sentences and entities to supplement the input textual context, serving as an ideal upper bound.
Appendix D Result of Textual Input Experiment
We present the full evaluation results in Table 4.
Appendix E Analysis of Entity Generation
@@
EAMA 50.64 13.16
Table 5: The Recall of Entities in and out of textual input context on NYTimes800k dataset. All methods except EAMA take loc_context as textual input, EAMA takes supplemented textual input. We cite the results of Tell and JoGANIC from previous researchesZhang and Wan (2023).
In-Out Textual Input Entity Generation According to prior researchZhang and Wan (2023), evaluating the generation of entities that appear in both the textual input context and captions can provide better insight into the models’ ability on handling entity information. Therefore, we further analyze the Recall of entities in and outside of the textual input context.
@@
All results are listed in Table 5, InstructBLIP obtain high Recall especially for entities of caption occurring in the textual input context. Our method obtains the highest Recall of entities appearing both in and outside of textual input context, which is consistency with the our previous results.
@@
EAMA 29.70 28.21
Table 6: The Precision and Recall of Entities which are out of train dataset in our approach: EAMA, on NYTimes800k dataset.
Out-of-Train Entity Generation To assess our method’s capability to predict entities not present in the training data, we examine captions in the NYTimes800k test dataset, focusing on entities that do not appear in the training data. The entire test set comprises 70,002 entities in the reference captions, with 13,481 of them not appearing in the textual input contexts or captions of the training data. Our approach, EAMA, successfully predicts 3,803 entities in the generated captions, resulting in a Recall of 28.21. The Precision of the out-of-train entities is 29.70, as shown in Table 6. Both the Recall and Precision metrics are comparable to the results for all entities presented in Section 5.2, demonstrating the generalization ability of EAMA for generating out-of-train entities.
Appendix F Case Study
We randomly select a sample from NYTimes800k dataset and generate captions using different models, as illustrated in Figure 5. From the generated captions, it’s evident that MLLMs under zero-shot settings can only produce common entity information such as "The United States" or "England", even when provided with the news image and associated text. Even the most powerful MLLM, GPT-4V (OpenAI et al., 2023) fails to generate specific entities related to the news image and mentioned in the news article, such as "Julie Ertz" and "Lucy Bronze." After OSFT on NYTimes800k dataset, InstructBLIP still cannot generate the names of people in the news image. LLaVA-v1.5 successfully generates "Bronze" but overlooks "England." In contrast, our method generates both common entities and the names of people mentioned in the news article, further affirming the effectiveness of our approach in entity recognition.
We also show another example to further illustrate the effectiveness of our method in figure 6. Our method achieves the best generation on entities and captions on both examples.
Figure 4: An example of our prompts for Entity-Aware Sentence Selection, Entity Selection and News Image Captioning tasks. We label entities in the caption in green. We label entities in the input context but not in the caption in blue.
Figure 5: An example of news image caption generation. We present the news image, news article and the news image caption of one sample from NYTimes800k dataset. The news image captions are respectively, generated in: (1) InstructBLIP in zero-shot setting. (2) LLaVA-v1.5 in zero-shot setting. (3) MiniGPT-v2 in zero-shot setting. (4) GPT-4V in zero-shot setting. (5) InstructBLIP (OSFT) on NYTimes800k. (6) LLaVA-v1.5 (OSFT) on NYTimes800k. (7) MiniGPT-v2 (OSFT) on NYTimes800k. (8) Our method, EAMA, on NYTimes800k. (9) is the ground-truth caption. We highlight Entities occurred in reference in red.
Figure 6: An example of news image caption generation. We present the news image, news article and the news image caption of one sample from NYTimes800k dataset. The news image captions are respectively, generated in: (1) InstructBLIP in zero-shot setting. (2) LLaVA-v1.5 in zero-shot setting. (3) MiniGPT-v2 in zero-shot setting. (4) GPT-4V in zero-shot setting. (5) InstructBLIP (OSFT) on NYTimes800k. (6) LLaVA-v1.5 (OSFT) on NYTimes800k. (7) MiniGPT-v2 (OSFT) on NYTimes800k. (8) Our method, EAMA, on NYTimes800k. (9) is the ground-truth caption. We highlight Entities occurred in reference in red.